# Calib3D: Calibrating Model Preferences for Reliable 3D Scene Understanding

### v1

@@ 5 Conclusion
In this study, we have presented Calib3D, a pioneering benchmark aimed at evaluating the reliability and accuracy of uncertainty estimates of 3D scene understanding models. Through our comprehensive assessment of state-of-the-art models across diverse 3D datasets, we have illuminated the critical challenges these models face in delivering confident and accurate predictions, especially within the context of safety-critical applications. Our findings reveal a significant gap in the calibration capabilities of current 3D scene understanding models, which often exhibit commendable accuracy but fall short in aligning their confidence levels with their predictive accuracy. Addressing this crucial gap, we introduced DeptS, a depth-aware scaling method specifically tailored to enhance the calibration of 3D models. By dynamically adjusting the logits distribution based on depth-correlated temperature scaling, the calibration performance is significantly improved. We hope that Calib3D and DeptS will inspire further research and innovation in the field of reliable 3D scene understanding.
@@ 6 Calib3D Benchmark
Table 8: Summary of 3D datasets encompassed in the Calib3D benchmark, including 1nuScenes [29], 2SemanticKITTI [7], 3Waymo Open [99], 4SemanticPOSS [83], 5SemanticSTF [115], 6ScribbleKITTI [107], 7Synth4D [95], 8S3DIS [3], and 9nuScenes-C and 10SemanticKITTI-C from the Robo3D benchmark [55]. Each dataset sheds light on a specific data acquisition and annotation protocol, such as different LiDAR sensor setups, adverse weather conditions, weak annotations, synthetic data, indoor scenes, and out-of-domain corruptions. The images are adopted from the original papers.
nuScenes SemanticKITTI Waymo Open SemanticPOSS SemanticSTF
ScribbleKITTI Synth4D S3DIS nuScenes-C SemanticKITTI-C
@@ 7.3 PyTorch-Style ECE Calculation
To facilitate reproduction, we provide a PyTorch-style code snippet for calculating the expected calibration error (ECE) on point clouds in LABEL:lst:ece3d.
7.4 PyTorch-Style Implementation of DeptS
To facilitate reproduction, we provide a PyTorch-style code snippet of the proposed depth-aware scaling (DeptS) method in LABEL:lst:depts.
@@ 8.2 Domain-Shift Uncertainty Estimation
Enhancing the uncertainty estimation capability of handling challenging scenarios is crucial for the practical usage of 3D scene understanding systems. We supplement the domain-shift uncertainty estimation results of FRNet [122] and SPVCNN [101] in Tab. 10 and Tab. 11, respectively. Similar to the observations drawn in the main body of this paper, we find that 3D models are vulnerable under adverse conditions. The expected calibration errors are extremely high under “fog”, “motion blur”, “crosstalk”, and “cross sensor” corruptions, which are commonly occurring scenarios in the real world. Compared to previous calibration methods like temperature scaling and meta-calibration, our DeptS shows a more stable performance across different domain-shift scenarios. We believe such an ability will become more and more important in the future development of 3D scene understanding systems.
Table 10: The expected calibration error (ECE) of FRNet [122] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [55]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
@@
Table 11: The expected calibration error (ECE) of SPVCNN [101] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [55]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
@@
Figure 6: The reliability diagrams of randomly sampled model predictions generated by CENet [16] on the validation set of the SemanticKITTI [7] dataset. UnCal, TempS, LogiS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, meta, and our proposed depth-aware scaling calibrations, respectively.

### v2

@@ 5 Conclusion
We introduced Calib3D, a benchmark that focuses on evaluating the reliability of uncertainty estimates in 3D scene understanding models. Through extensive evaluations of state-of-the-art models across diverse 3D datasets, we highlighted critical challenges in delivering confident and accurate predictions, particularly in safety-critical applications. Our results expose a significant gap in the calibration of current 3D models, which often achieve high accuracy but struggle to align confidence with predictive accuracy. To address this, we proposed DeptS, a depth-aware scaling method that enhances calibration by adjusting logits based on depth-correlated temperature scaling. We hope that Calib3D and DeptS will inspire further research and innovation in the field of reliable 3D scene understanding.
Acknowledgments. This work is supported by the Ministry of Education, Singapore, under MOE AcRF Tier 2 (MOET2EP20221-0012), NTU NAP, and RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ 6 Calib3D Benchmark
Table 8: Summary of 3D datasets encompassed in the Calib3D benchmark. A total of ten 3D datasets have been used in our benchmark, including 1nuScenes [31], 2SemanticKITTI [7], 3Waymo Open [108], 4SemanticPOSS [92], 5SemanticSTF [125], 6ScribbleKITTI [116], 7Synth4D [104], 8S3DIS [3], and 9nuScenes-C and 10SemanticKITTI-C from the Robo3D benchmark [59]. Each dataset sheds light on a specific data acquisition and annotation protocol, such as different LiDAR sensor setups, adverse weather conditions, weak annotations, synthetic data, indoor scenes, and out-of-domain corruptions. The images shown here are adopted from the original dataset papers.
nuScenes SemanticKITTI Waymo Open SemanticPOSS SemanticSTF
ScribbleKITTI Synth4D S3DIS nuScenes-C SemanticKITTI-C
@@ 7.3 PyTorch-Style ECE Calculation
To facilitate reproduction, we provide a PyTorch-style code snippet for calculating the expected calibration error (ECE) on point clouds as follows.
@@
7.4 PyTorch-Style Implementation of DeptS
To facilitate reproduction, we provide a PyTorch-style code snippet of the proposed depth-aware scaling (DeptS) method as follows.
@@ 8 Additional Quantitative Result
8.1 Depth Correlations in LiDAR Data
As discussed in Sec. 3.3 of the main body of this paper, the motivation behind the depth-aware scaling method, DeptS, stems from some interesting observations from our experiments. We observe that traditional calibration techniques, effective in 2D image-based tasks, struggle with 3D data due to the unique characteristics of point clouds, such as being unordered and lacking texture. Through our analysis, we identified a clear correlation between calibration errors, prediction entropy, and depth. Specifically, as shown in Fig. 5, LiDAR points at greater distances from the ego vehicle often exhibit lower accuracy, yet uncalibrated models maintain high confidence in these areas, leading to substantial calibration errors [62, 64]. This overconfidence in distant regions prompted the need for a tailored approach to address the depth-related calibration issue.
To tackle this, we propose DeptS, a method that adjusts model confidence based on depth information. By introducing a depth-correlation coefficient that reweights the temperature scaling parameters, DeptS reduces confidence for LiDAR points at larger depths, effectively mitigating the overconfidence problem. This method allows for better calibration in 3D scene understanding models, particularly in middle-to-far regions where predictions are less reliable, leading to improved calibration performance across diverse 3D datasets.
@@ 8.3 Domain-Shift Uncertainty Estimation
Enhancing the uncertainty estimation capability of handling challenging scenarios is crucial for the practical usage of 3D scene understanding systems [127, 42, 72, 59, 63, 16, 15, 60]. We supplement the domain-shift uncertainty estimation results of FRNet [133] and SPVCNN [110] in Tab. 11 and Tab. 12, respectively. Similar to the observations drawn in the main body of this paper, we find that 3D models are vulnerable under adverse conditions. The expected calibration errors are extremely high under “fog”, “motion blur”, “crosstalk”, and “cross sensor” corruptions, which are commonly occurring scenarios in the real world. Compared to previous calibration methods like temperature scaling and meta-calibration, our DeptS shows a more stable performance across different domain-shift scenarios. We believe such an ability will become more and more important in the future development of 3D scene understanding systems.
Table 10: Comparisons between the proposed DeptS and state-of-the-art network calibration methods on the validation set of SemanticKITTI [7]. All ECE (the lower the better) and mIoU (the higher the better) scores reported are in percentage (%).
Method Venue MinkUNet [20] CENet [18]
ECE mIoU ECE mIoU
UnCal - $3.04\%$ $63.05\%$ $5.95\%$ $60.87\%$
TempS [38] ICML’17 $3.01\%$ $63.05\%$ $3.93\%$ $60.87\%$
LogiS [38] ICML’17 $3.08\%$ $63.11\%$ $3.79\%$ $60.95\%$
MetaC [80] ICML’21 $2.69\%$ $62.93\%$ $3.31\%$ $60.81\%$
DeepEnsemble [69] NeurIPS’17 $2.99\%$ $64.95\%$ $5.61\%$ $61.70\%$
BatchEnsemble [121] ICLR’20 $2.77\%$ $64.70\%$ $5.40\%$ $62.13\%$
MIMO [43] ICLR’21 $3.21\%$ $63.60\%$ $6.97\%$ $61.62\%$
PackedEnsemble [70] ICLR’23 $2.82\%$ $63.88\%$ $6.00\%$ $59.81\%$
DeptS Ours $\mathbf{2.63\%}$ $\mathbf{63.47\%}$ $\mathbf{3.09\%}$ $\mathbf{61.20\%}$
Table 11: The expected calibration error (ECE, the lower the better) of FRNet [133] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [59]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibration methods, respectively.
@@
Table 12: The expected calibration error (ECE, the lower the better) of SPVCNN [110] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [59]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibration methods, respectively.
@@
8.4 Comparisons to Recent Calibration Methods
In the main body of this paper, we provide a comprehensive benchmark study of classical network calibration methods, such as TempS, LogiS, DiriS, and MetaC, across a range of ten different 3D datasets. The benchmark results verify that the proposed DepthS exhibits stronger performance compared to these classical approaches.
To provide a more holistic evaluation of DepthS compared to more recent network calibration methods, we conduct experiments with more recent network calibration methods, including DeepEnsemble [69], BatchEnsemble [121], MIMO [43], and PackedEnsemble [70], on the validation set of the SemanticKITTI [7] dataset. As shown in Tab. 10, the results demonstrate that our proposed DeptS is consistently better than both the classical and recent network calibration methods.
Figure 6: The reliability diagrams of randomly sampled model predictions generated by the CENet [18] model on the validation set of the SemanticKITTI [7] dataset. UnCal, TempS, LogiS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, meta, and our proposed depth-aware scaling calibration methods, respectively.